# Radial Gradients and Intrinsic Scatter in MaNGA Galaxies

### v1

@@ Abstract
We derive stellar population parameters and their radial gradients within 0.65 Re for spatially resolved spectra of 2417 early-type galaxies from the MaNGA survey with stellar velocity dispersions ($\sigma$) between 50 kms-1 and 340 kms-1. We invert a grid of metallicity-composite stellar population models to find mean age and the abundances of C, N, Na, Mg, and Fe. These models have significant improvements compared to past models, including isochrones that respond to individual element abundances.
Globally, age rises with $\sigma$ while [Fe/H] gently falls. Individual light element abundances strengthen with $\sigma$ but strengthen faster for log($\sigma$)$>$2.0. [Fe/H] shows a maximum at log($\sigma$$\approx$2.0, falling to either side. Light element [X/Fe] anticorrelate with [Fe/H]. Heterogeneity as measured by astrophysical scatter is highest in low-$\sigma$ galaxies, most dramatically for age, Fe, and N.
For galaxy-internal parameters, age shows nearly flat radial gradients in low-$\sigma$ galaxies, slightly negative at high $\sigma$. The mean radial gradient in [Fe/H] is negative and light element [X/Fe]s fall. Intrinsic scatter in gradients is highest in high-$\sigma$ galaxies, most dramatically for age and Fe. Evidently, nearly as many galaxies form inside-out as form outside-in.
A near-zero radial gradient in age and light elements coupled with a mild [Fe/H] gradient supports the hierarchical merging scenario for ETG evolution. IllustrisTNG hierarchical simulations reproduce the age structure we find, show the abundance slope changes at log($\sigma$$\approx$2.0 that we observe, and exhibit flat gradients similar to those we derive, although the abundances predicted by IllustrisTNG are significantly higher than our observations overall.
@@ 1 Introduction
The evolution of the chemical properties of stellar populations, as well as the interstellar and intergalactic medium across cosmic epochs, offers valuable insights into the processes that drive galaxy formation and evolution. A wide range of models and cosmological simulations have been created to explore the evolution of the baryonic matter, star formation processes, and the characteristics of galaxies throughout cosmic history, including in the local universe (Springel et al., 2005; Vogelsberger et al., 2014a; Somerville & Davé, 2015; Schaye et al., 2015; Nelson et al., 2019). The abundances of certain heavy elements relative to hydrogen put strong constraints on these models. The metal content of galaxies, often tracked via [Fe/H], reflects heavy elements added by supernovae and other stellar sources such as nucleosynthetically-enriched mass loss and neutron star mergers minus the amount of metals removed through mechanisms such as outflows and stripping (Tremonti et al., 2004; Finlator & Davé, 2008; Peeples & Shankar, 2011; De Rossi et al., 2015; Chisholm et al., 2017). Heavy element abundances are also affected by dilution caused by inflows of pristine gas. Iron is a key element in understanding the evolution of galaxies (Kereš et al., 2005; Davé et al., 2011; Sánchez Almeida et al., 2014) but various other chemical elements are formed over different timescales by different types of stars. The hope is that their relative abundances can be translated into the star formation history and chemical evolution within a galaxy (Tinsley, 1979; Matteucci, 1996; Thomas et al., 2003; Nomoto et al., 2013; Weinberg et al., 2017).
@@
Analyzing [Fe/H] and chemical abundances on spatially resolved scales, often summarized as radial gradients, may provide further insight into the processes that have shaped the growth and assembly of galaxies (Zaritsky et al., 1994; Rawle et al., 2010; Li et al., 2013; Goddard et al., 2017; La Barbera et al., 2017; Esteban et al., 2020). Studies on [Fe/H] and elemental abundance gradients also offer information on whether galaxies formed “inside-out” or “outside-in” (Chiappini et al., 1997; Garnett, 2004; Kobayashi & Nakasato, 2011), and elucidate the role of internal processes such as galactic fountains, stellar migration, and radial gas inflows (Dekel & Silk, 1986; Sellwood & Binney, 2002; Oppenheimer & Davé, 2008; Roškar et al., 2008b, a; Spitoni et al., 2009; Melioli et al., 2015).
@@
The formation history and evolution of galaxies remain a subject of considerable debate. Two prominent scenarios provide convenient bookends for discussion. The monolithic collapse hypothesis suggests that galaxies form from the dissipation of large, pre-galactic gas clouds, leading to the formation of all their stars in one fairly brief event (Larson, 1974, 1976; Thomas et al., 1999; Pipino et al., 2010). According to this model, galaxies are predicted to form via “outside-in” formation, where star formation in the outermost regions ceases earlier due to the earlier onset of galactic winds in areas with shallower gravitational potential wells (Pipino et al., 2006). Conversely, the hierarchical merger hypothesis proposes that large galaxies form through the merging of smaller subunits (Kauffmann et al., 1993; Kauffmann & Charlot, 1998; De Lucia & Blaizot, 2007). In this scenario, star formation is an ongoing process that depends on the gas content of the progenitor galaxies. Galaxies formed through the merging of smaller units tend to have flattened [Fe/H] gradients (White, 1980; Di Matteo et al., 2009). In contrast, galaxies formed through dissipative processes are expected to have steep to moderate gradients in [Fe/H] and other heavy element abundances (Larson, 1976; Carlberg & Freedman, 1985). There have also been studies where the fundamentals of both of these theories are shown to be partially true (Mehlert et al., 2003; Ogando et al., 2005; Goddard et al., 2017). An overview of both of these theories can be found in Sil’chenko (2015).
@@
Stellar population analysis works at all redshifts. It is an important tool in analyzing the chemical compositions of the galaxies that hold clues to the galaxy formation and evolution over many epochs of time (Walcher et al., 2011; Conroy, 2013). It is possible that formation time-scales in galaxies can be measured via the ratio of $\alpha$-elements to that of heavy metals (Fe-peak elements) (Thomas et al., 1999), although the high-mass initial mass function (IMF) may potentially also alter it (Worthey et al., 1992, 2014). Abundances of other lighter metals like carbon (C) and nitrogen (N) put constraints on star formation timescale and also gas inflow (Chiappini et al., 2003; Johansson et al., 2012; Berg et al., 2019). Spectral analysis of galaxies have also been shown to be effective in constraining the slope in the low-mass IMF of galaxies (Martín-Navarro et al., 2015; van Dokkum et al., 2017; Parikh et al., 2018).
@@
The majority of existing studies has concentrated on detailed stellar population analysis of Early-Type Galaxies (ETGs). This focus is due to the fact that ETGs can be more accurately modeled using simple star formation histories, as they typically exhibit no ongoing star formation (Bruzual & Charlot, 2003; Vazdekis et al., 2010; Worthey et al., 2014) and thus allow analysis of Balmer features that are otherwise susceptible to emission line fill-in. ETGs are predominantly composed of older stars, with minimal or no current star formation activity. In favorable cases, at least, they have had only modest changes over much of cosmic time, and present themselves as good laboratories for studying the long-term evolution of stellar populations.
@@
The study of gradients in various parameters has long been used as a mean to understand the formation and evolution of galaxies. Two of the first studies available in literature on color gradients were done by Tinsley & Larson (1977) and Franx & Illingworth (1990). In the early 1990s, a number of pioneering studies were published on line strength gradients, by Peletier (1989); Gorgas et al. (1990), Bender & Surma (1992), Davies et al. (1993), and others. The first mention of an [Fe/H] gradient (as opposed to a line strength gradient unattached to the solar-zeroed [Fe/H] standard) that we found in literature is by Carollo et al. (1993) where they measured line strengths of different features from 42 ETGs to infer an [Fe/H] gradient. These studies did not necessarily use any stellar population model to infer age, [Fe/H], or element abundances but they pioneered the field of gradient studies. Later on with the development reliable stellar population models, a large number of studies had been completed on spatial gradients in age, [Fe/H], and element abundances within both ETGs and Late Type Galaxies (LTGs) (Mehlert et al., 2003; Forbes et al., 2005; Proctor et al., 2005; Jablonka et al., 2007; Spolaor et al., 2009; Rawle et al., 2010; Kuntschner et al., 2010; González Delgado et al., 2017; Goddard et al., 2017; Parikh et al., 2018; Feldmeier-Krause et al., 2021; Parikh et al., 2021). These studies generally agree that ETGs and the bulges of spiral galaxies display negative [Fe/H] gradients, while showing small positive or flat gradients in both age and the $\alpha$/Fe ratio. This “globalness” of age and $\alpha$ throughout the body of the galaxy is an example of a useful constraint to be matched by galaxy formation models.
@@
With the emergence of Integral Field Unit (IFU) surveys for galaxies opens up the scope of using high quality data to investigate gradients in stellar population parameters further. Such surveys include the SAURON and ATLAS-3D project (Bacon et al., 2001; Emsellem et al., 2004; Cappellari et al., 2011), CALIFA project (Sánchez et al., 2012), SAMI project (Croom et al., 2012), and MaNGA project (Bundy et al., 2015). IFU spectroscopy allows us to get galaxy spectra from different parts of the galaxy enabling detailed radial studies. Although these surveys provide us with a large volume of data, most of the studies till date are either limited by sample size (highest number of individual galaxies studied till date is by Goddard et al. (2017) where they studied around 700 ETGs and Ferreras et al. (2019) analyzing around 500 ETGs) or use the spectra stacking technique (Parikh et al., 2018) which when used looses the individual information of a galaxy. Some other studies also treat each galaxy individually but they are severely restricted in terms of sample size (Greene et al., 2015; Alton et al., 2018; Feldmeier-Krause et al., 2021). In this work, we consider 2417 individual unique galaxies from the Mapping Nearby Galaxies at Apache Point Observatory (MaNGA, Bundy et al. (2015)) for detailed study on the age, [Fe/H], and element abundances along with their corresponding gradients.
The stellar population models employed are based on the framework established by Worthey (1994), incorporating the index definitions expanded by Serven et al. (2005), and further refined following Worthey et al. (2011). One of our focuses is on the age, [Fe/H], and C, N, sodium (Na), and magnesium (Mg) abundances for the galaxies and their dependence on stellar velocity dispersion ($\sigma$) and [Fe/H] (for abundances only). We take a deep dive into the radial gradients in those parameters and how they are changing with $\sigma$. Alongside these, we also calculate the intrinsic scatter ($\delta_{i}$) in all the measured quantities and discuss the results.
@@
It is important to state at the outset a couple of important recent improvements in the models that separate these results from previous work. First, the isochrones upon which the models rest now vary with individual element abundances (Worthey et al., 2022). This enables an effect we term “magnesium amplification” that applies to Mg and other elements as well. Addition of Mg, alone, to the chemical mixture in a stellar population, will add opacity to the stars, increase their radii, and cool them. Cooler stars have stronger Mg lines. So, not only will Mg line strengths increase due to the increased number of absorbers, Mg lines will also increase in strength due to arising in cooler stars. Since inclusion of this effect is new to the field starting with this work, our results will not necessarily agree with previous estimates of quantities such as [Fe/H], [C/Fe], or [Mg/Fe]. Second, we incorporate an abundance distribution function (ADF) instead of a delta function in abundance (Tang et al., 2014). It is important to do this because metal-poor populations are brighter in the blue, and thus dilute metallic lines found in that spectral region. In the red, the balancing favors metal-rich populations more. We parameterize by the peak (mode) of the ADF, not by, for example, the $V$-weighted average abundance, or, for example, the mass-weighted average abundance. This introduces modest, well-behaved offsets between results derived from a simple stellar population (SSP) and this work, as illustrated in Tang et al.
The influence of uncertainties in stellar population modelling has been well documented by Conroy et al. (2009). A lot of fundamental work was also done in 1990s on the effect of intrinsic scatter in ETGs (Gregg, 1992; Guzman et al., 1993). A lot of references regarding the effect of the intrinsic scatter on the Fundamental Plane (FP) relationship can be found in literature (Forbes et al., 1998; Blakeslee et al., 2009; Hyde & Bernardi, 2009; Magoulas et al., 2012; D’Eugenio et al., 2024). By including [$\alpha$/Fe] as a fourth parameter in the FP, Gargiulo et al. (2009) showed that the total scatter reduces from 0.088 to 0.075 dex, while the estimated intrinsic scatter decreases from 0.068 to 0.049 dex. Falcón-Barroso et al. (2011) quantified intrinsic scatter in scaling relations for nearby elliptical galaxies.
@@
However, the intrinsic, astrophysical scatter in age, [Fe/H], and elemental abundances are not well documented in previous literature. We wanted to better this situation and thus treated each MaNGA galaxy in our study individually (not using any kind of stacking algorithm).
@@
This article is arranged as follows: §$\S$2 explains the source and type of data used in this work, §$\S$3 contains the details of the analysis performed along with brief explanation of the stellar population model used for this work, §$\S$4 elucidates all of our results, §$\S$5 discusses the implications of our results and comparison with previous works, and also discusses our results on the intrinsic scatter. §$\S$6 concludes the article with reiterating our main results and forecasting future work, and an appendix gives a library of graphical comparisons with the TNG-100 cosmological evolution simulation, part of tables elucidating all the indices used in this work and result of the inversion program compfit.

### v2

@@ Abstract
We derive stellar population parameters and radial gradients within 0.65 Re for spatially resolved spectra of 2968 early-type galaxies from MaNGA, spanning stellar velocity dispersions ($\sigma$) of 50–340 km s-1. Light weighted mean age, C, N, Na, Mg and Fe abundances are obtained by inverting metallicity–composite stellar population models with isochrones that respond in $T{\rm{}_{eff}}$ to individual element abundance changes.
Globally, log age increases ($\sim$0.53 dex per decade), [Fe/H] declines slightly ($\sim$–0.06 dex per decade), and [X/Fe] for light elements rises (0.19–0.37 dex per decade for log $\sigma<$ 2.0 but steepens to nearly double slope for log $\sigma>$ 2.0) with log $\sigma$. [Fe/H] peaks at log $\sigma\approx$2.0 and falls on either side. Light element [X/Fe] anticorrelates with [Fe/H] ($\sim$–0.1 dex per decade). Astrophysical scatter is largest in low-$\sigma$ galaxies, especially for Fe and N.
Internally, age gradients are nearly flat in low-$\sigma$ galaxies and slightly negative in high-$\sigma$ systems ($\sim$–0.04 dex per decade). [Fe/H] radial gradients steepen from –0.06 dex per decade to –0.15 dex per decade across $\sigma$, while light elements (except Na) show $\sim$–0.03 dex per decade gradients. Scatter in gradients peaks in high-$\sigma$ galaxies, most strongly for Fe ($\sim$0.23 dex), suggesting comparable numbers of inside-out and outside-in formation.
A near-zero ($\sim$–0.03) age and light-element gradient plus mild [Fe/H] gradients supports hierarchical merging for ETG evolution. Simulations match the observed age structure, slope change at log $\sigma\approx$2.0, and flat gradients, though they overpredict absolute abundances.
@@ 1.1 Galaxy Evolution
The cosmic evolution of the chemical properties of stellar populations, along with that of the interstellar and intergalactic medium, offers valuable insights into the processes that drive galaxy formation and evolution. A wide range of models and cosmological simulations have been created to explore the evolution of the baryonic matter, star formation processes, and the characteristics of galaxies throughout cosmic history, including in the local universe (Springel et al., 2005; Vogelsberger et al., 2014a; Somerville & Davé, 2015; Schaye et al., 2015; Nelson et al., 2019). The abundances of certain heavy elements relative to hydrogen constrain these models. The metal content of galaxies, often traced via [Fe/H], reflects heavy elements added by supernovae and other stellar sources such as nucleosynthetically-enriched mass loss and neutron star mergers, minus the amount of metals removed through mechanisms such as outflows and stripping (Tremonti et al., 2004; Finlator & Davé, 2008; Peeples & Shankar, 2011; De Rossi et al., 2015; Chisholm et al., 2017). Heavy element abundances are also affected by dilution caused by inflows of pristine gas. Iron is a key element in understanding the evolution of galaxies (Kereš et al., 2005; Davé et al., 2011; Sánchez Almeida et al., 2014) but various other chemical elements are formed over different timescales by different types of stars, and these elements might also offer insights. The hope is that their relative abundances can be translated into the star formation history and chemical evolution within a galaxy (Tinsley, 1979; Matteucci, 1996; Thomas et al., 2003; Nomoto et al., 2013; Weinberg et al., 2017).
@@
Analyzing chemical abundances on spatially resolved scales, often summarized as radial gradients, may provide further insight into the processes that have shaped the growth and assembly of galaxies (Zaritsky et al., 1994; Rawle et al., 2010; Li et al., 2013; Goddard et al., 2017; La Barbera et al., 2017; Esteban et al., 2020). Studies on [Fe/H] and elemental abundance gradients also offer information on whether galaxies formed “inside-out” or “outside-in” (Chiappini et al., 1997; Garnett, 2004; Kobayashi & Nakasato, 2011; Sil’chenko, 2015), and elucidate the role of internal processes such as galactic fountains, stellar migration, and radial gas inflows (Dekel & Silk, 1986; Sellwood & Binney, 2002; Oppenheimer & Davé, 2008; Roškar et al., 2008b, a; Spitoni et al., 2009; Melioli et al., 2015).
@@
The formation history and evolution of galaxies remain a subject of considerable debate. Two prominent scenarios provide convenient bookends for discussion. The monolithic collapse hypothesis suggests that galaxies form from the dissipation of large, pre-galactic gas clouds, leading to the formation of all their stars in one fairly brief event (Larson, 1974, 1976; Thomas et al., 1999; Pipino et al., 2010). According to this model, galaxies are predicted to form via “outside-in” formation, where star formation in the outermost regions ceases earlier due to the earlier onset of galactic winds in areas with shallower gravitational potential wells and have strong metallicity gradients (Pipino et al., 2006). Conversely, the hierarchical merger hypothesis proposes that large galaxies form through the merging of smaller subunits (Kauffmann et al., 1993; Kauffmann & Charlot, 1998; De Lucia & Blaizot, 2007). In this scenario, star formation is an ongoing process that depends on the gas content of the progenitor galaxies. Galaxies formed through the merging of smaller units tend to flatten existing [Fe/H] gradients (White, 1980; Di Matteo et al., 2009). In contrast, galaxies formed through dissipative processes are expected to have steep to moderate gradients in [Fe/H] and other heavy element abundances (Larson, 1976; Carlberg & Freedman, 1985). There have also been studies where the fundamentals of both of these theories are shown to be partially true (Mehlert et al., 2003; Ogando et al., 2005; Goddard et al., 2017). An overview of both of these theories can be found in Sil’chenko (2015).
1.2 Models for the Integrated Light of Stellar Populations
@@
Stellar population analysis works at all redshifts. It is an important tool in analyzing the chemical compositions of the galaxies that hold clues to the galaxy formation and evolution over many epochs of time (Walcher et al., 2011; Conroy, 2013). It is possible that formation time-scales in galaxies can be measured via the ratio of $\alpha$-elements to that of heavy metals (Fe-peak elements) (Thomas et al., 1999), although the high-mass initial mass function (IMF) may potentially also alter it (Worthey et al., 1992, 2014b). Abundances of other lighter metals like carbon (C) and nitrogen (N) put constraints on star formation timescale and also gas inflow (Chiappini et al., 2003; Johansson et al., 2012; Berg et al., 2019). Spectral analysis of galaxies have also been shown to be effective in constraining the slope in the low-mass IMF of galaxies (Martín-Navarro et al., 2015; van Dokkum et al., 2017; Parikh et al., 2018).
@@
The majority of existing studies has concentrated on detailed stellar population analysis of Early-Type Galaxies (ETGs). This focus is due to the fact that ETGs can be more accurately modeled using simple star formation histories, as they typically exhibit no ongoing star formation (Bruzual & Charlot, 2003; Vazdekis et al., 2010; Worthey et al., 2014b) and thus allow analysis of Balmer features that are otherwise susceptible to emission line fill-in. ETGs are predominantly composed of older stars, with minimal or no current star formation activity. In favorable cases, at least, they have had only modest changes over much of cosmic time, and present themselves as good laboratories for studying the long-term evolution of stellar populations.
1.3 A Focus on Stellar Population Gradients
The first science focus for this paper is on the light-weighted age, [Fe/H], and C, N, sodium (Na), and magnesium (Mg) abundances for the galaxies and their dependence on stellar velocity dispersion ($\sigma$) and [Fe/H]. We also explore radial gradients in the same list of parameters and their $\sigma$ dependence. Finally, we compute the intrinsic scatter ($\delta_{i}$) in all the quantities and discuss the trends.
@@ 1.3 A Focus on Stellar Population Gradients
The study of gradients in various parameters has long been used as a mean to understand the formation and evolution of galaxies. Two of the first studies available in literature on color gradients were done by Tinsley & Larson (1977) and Franx & Illingworth (1990). In the early 1990s, a number of pioneering studies were published on line strength gradients, by Peletier (1989); Gorgas et al. (1990), Bender & Surma (1992), Davies et al. (1993), and others. The first mention of an [Fe/H] gradient (as opposed to a line strength gradient unattached to the solar-zeroed [Fe/H] standard) that we found in literature is by Carollo et al. (1993) where they measured line strengths of different features from 42 ETGs to infer an [Fe/H] gradient. These studies did not necessarily use any stellar population model to infer age, [Fe/H], or element abundances but they pioneered the field of gradient studies. Later on with the development reliable stellar population models, a large number of studies had been completed on spatial gradients in age, [Fe/H], and element abundances within both ETGs and Late Type Galaxies (LTGs) (Mehlert et al., 2003; Forbes et al., 2005; Proctor et al., 2005; Jablonka et al., 2007; Spolaor et al., 2009; Rawle et al., 2010; Kuntschner et al., 2010; González Delgado et al., 2017; Goddard et al., 2017; Parikh et al., 2018; Li et al., 2018; Ferreras et al., 2019; Domínguez Sánchez et al., 2020; Feldmeier-Krause et al., 2021; Parikh et al., 2021; Lu et al., 2023). These studies generally agree that ETGs and the bulges of spiral galaxies display negative [Fe/H] gradients, while showing small positive or flat gradients in both age and the $\alpha$/Fe ratio. This “globalness” of age and $\alpha$ throughout the body of the galaxy is an example of a useful constraint to be matched by galaxy formation models.
@@
The emergence of Integral Field Unit (IFU) surveys for galaxies opens up the scope of using high-quality data to investigate gradients in stellar population parameters further. Such surveys include the SAURON and ATLAS-3D project (Bacon et al., 2001; Emsellem et al., 2004; Cappellari et al., 2011), CALIFA project (Sánchez et al., 2012), SAMI project (Croom et al., 2012), and MaNGA project (Bundy et al., 2015). IFU spectroscopy allows us to get galaxy spectra from different parts of the galaxy enabling detailed radial studies. Although these surveys provide us with a large volume of data, most of the studies to date are either limited by sample size (the largest number of individual galaxies studied to date is by Goddard et al. (2017) where they studied around 700 ETGs and Ferreras et al. (2019) at about 500 ETGs) or the studies stack spectra (Parikh et al., 2018). Stacking suppresses information on individual galaxies. Other studies treat each galaxy individually but are restricted in sample size (Greene et al., 2015; Alton et al., 2018; Feldmeier-Krause et al., 2021). In this work, we consider 2968 individual unique galaxies from the Mapping Nearby Galaxies at Apache Point Observatory (MaNGA, Bundy et al. (2015)) for a detailed study of the age, [Fe/H], and element abundances along with their corresponding gradients.
1.4 A Focus on Exploiting Improved Models
@@
It is important to state at the outset a couple of important recent improvements in the models that separate these results from previous work. First, the isochrones upon which the models rest now vary with individual element abundances (Worthey et al., 2022). This enables an effect we term “magnesium amplification” that applies to Mg and other elements as well. Addition of Mg, alone, to the chemical mixture in a stellar population, will add opacity to the stars, increase their radii, and cool them. Cooler stars have stronger Mg lines. So, not only will Mg line strengths increase due to the increased number of absorbers, Mg lines will also increase in strength due to arising in cooler stars. Since inclusion of this effect is new to the field starting with this work, our results will not necessarily agree with previous estimates of quantities such as [Fe/H], [C/Fe], or [Mg/Fe]. Second, we incorporate an abundance distribution function (ADF) instead of a delta function in abundance (Tang et al., 2014). It is important to do this because metal-poor populations are brighter in the blue, and thus dilute metallic lines found in that spectral region. In the red, the balancing favors metal-rich populations more. We parameterize by the peak (mode) of the ADF, not by, for example, the $V$-weighted average abundance, or, for example, the mass-weighted average abundance. This introduces modest, well-behaved offsets between results derived from a simple stellar population (SSP) and this work, as illustrated in Tang et al. (2014).
1.5 A Focus on Astrophysical Scatter
The influence of uncertainties in different phases of stellar evolution and IMF on derived stellar population parameters like the stellar masses, mean ages, metallicities, star formation histories, and other has been well documented by Conroy et al. (2009) (see also Conroy et al. (2010) and Conroy & Gunn (2010)). A lot of fundamental work was also done in 1990s on the effect of intrinsic scatter in ETGs (Gregg, 1992; Guzman et al., 1993). Falcón-Barroso et al. (2011) quantified intrinsic scatter in scaling relations for nearby elliptical galaxies.
@@ 1.5 A Focus on Astrophysical Scatter
However, intrinsic, astrophysical scatter in age, [Fe/H], and elemental abundances goes unstudied in recent literature. To improve this situation, we treated each MaNGA galaxy in our study individually (not using any kind of stacking algorithm).
@@
This article is arranged as follows: $\S$2 explains the source and type of data used in this work, $\S$3 contains the details of the analysis performed along with brief explanation of the stellar population model used for this work, $\S$4 elucidates all of our results, $\S$5 discusses the implications of our results and comparison with previous works, and also discusses our results on the intrinsic scatter. $\S$6 concludes the article with reiterating our main results and forecasting future work, and an appendix gives a library of graphical comparisons with the TNG-100 cosmological evolution simulation, part of tables elucidating all the indices used in this work and result of the inversion program compfit.